Hart Labs, Inc. provides mad cow disease testing for both state and federal governmental agricultural agencies. Because the company’s customers are governmental agencies, prices are strictly regulated. Therefore, Hart Labs must constantly monitor and control its testing costs. Shown below are the standard costs for a typical test.

Direct materials (2 test tubes @ $1.40 per tube) $2.80
Direct labor (1 hour @ $32 per hour) 32.00
Variable overhead (1 hour @ $6.00 per hour) 6.00
Fixed overhead (1 hour @ $12.00 per hour) 12.00
    Total standard cost per test $52.80


The lab does not maintain an inventory of test tubes. As a result, the tubes purchased each month are used that month. Actual activity for the month of November 2020, when 1,400 tests were conducted, resulted in the following.

Direct materials (2,940 test tubes) $3,822
Direct labor (1,456 hours) 45,136
Variable overhead 8,316
Fixed overhead 15,792


Monthly budgeted fixed overhead is $17,520. Revenues for the month were $91,000, and selling and administrative expenses were $3,600.

(a)

Compute the price and quantity variances for direct materials and direct labor.

Materials price variance $                                                                          Neither favorable nor unfavorableFavorableUnfavorable
Materials quantity variance $                                                                       Neither favorable nor unfavorableUnfavorableFavorable
Labor price variance $                                                                       UnfavorableNeither favorable nor unfavorableFavorable
Labor quantity variance $                                                                       FavorableNeither favorable nor unfavorableUnfavorable

b) Compute the total overhead variance

Total Overhead Variance: $

C) Prepare an income statement for management.
Answer #1

answer 1.

MATERIAL PRICE VARIANCE $ 294 FAVORABLE (AP < SP )
MATERIAL QUANTITY VARIANCE 205.8 UNFAVORABLE ( AQ>SQ)
LABOR PRICE VARIANCE 1456 FAVORABLE ( AP < SP )
LABOR EFFICIENCY 1792 UN FAVORABLE ( AH>SH)

A)

price variance = (AP - SP )* AQ

actual price = 1.3 per test tube

3822 / 2940 = 1.3

standard price = 1.4

actual quantity = 2940

(1.4 - 1.3 ) * 2940

=294

material price variance = 294 it is favorable ( actual price < standard price )

b )

material quantity variance = (AQ - SQ ) SP

AQ = 2947

SQ = 2800

(2 * 1400 UNIT = 2800)

SP = 1.40

( 2947 - 2800 ) * 1.40

= 205.8

material quantity variance = 205.8 UNFAVORABLE ( ACTUAL QUANTITY > STANDARD QUANTITY )

C )

LABOR PRICE VARIANCE = ( AP - SP ) * AQ

AP = 31

45136 / 1456 = 31

SP = 32

AQ = 1456 HR

( 32 - 31 ) 1456

LABOR PRICE VARIANCE = 1456 IT IS FAVORABLE ( AP < SP )

D )

LABOR QUANTITY VARIANCE = ( AH - SH ) SP

ACTUAL HOUR = 1456

STANDARD HOUR = 1400

(1 * 1400 )

STANDARD PRICE PER HOUR = 32

( 1456 - 1400 ) 32

LABOR QUANTITY VARIANCE = 1792 IT IS UNFAVORABLE ( ACTUAL HOUR > STANDARD HOUR )

ANSWER 2.

TOTAL OVER OVERHEAD VARIANCE

Overhead applied = Overhead application rate x SH

Overhead variance = Overhead applied – Actual overhead

ACTUAL OVERHEAD = 24108

FIXED OVERHEAD = 15792

VARIABLE OVERHEAD = 8316

actual over head = 25920

standard fixed overhead = 17520

standard variable overhead = 8400

applied overhead = 25920

25920 - 24108 = 1812 favorable ( actual overhead < standard overhead)

answer 3.

income statement

particular amount amount
sales 91,000
cost of goods sold ( DM+ DL + V O/H) 57274
manufacturing margin 33726
selling and administration 3600
contribution 30126
fixed cost 15792
operating income 14334